OpenStudy (imtiaz7):

Find the LCM of the following expressions by factorization. 2(x^4-y^4) , 3(x^3+2x^2*y-xy^2-2y^3)

OpenStudy (zzr0ck3r):

can you factor the left?

OpenStudy (zzr0ck3r):

2(x^4-y^4)= 2(x^2-y^2)(x^2+y^2) = 1(x-y)(x+y)(x^2+y^2)

OpenStudy (anonymous):

2(x+y)(x-y)(x^2+y^2) and 3(x^2(x+2y)-y^2(x+2y))=3(x+y)(x-y)(x+2y) hence lcm is 6(x+y)(x-y)(x^2+y^2)(x+2y)

OpenStudy (imtiaz7):

guys the answer in book is: 6(x+2y)(x^4-y^4)

OpenStudy (anonymous):

thats what i wrote 6 (x^2-y^2)(x^2+y^2)(x+2y) = 6 (x+2y)(x^4-y^4 )

OpenStudy (zzr0ck3r):

the book is not simplifying as far as possible